please help me answer the following equation:
1.) rearrange the following equation to find μ:
0.6=e^μ*2.1

Answer by Theo(13342)

let m = mu.

.6 = e^(m * 2.1)

take the natural log of both sides of the equation to get:

ln(.6)) = ln(e^(m * 2.1))

since ln(e^x) = x * ln(e) and since ln(e) = 1, the equation becomes:

ln(.6) = m * 2.1

solve for m to get m = ln(.6) / 2.1 = -.243250297

replace m in the original equation with this value to see if the original equation holds true.

original equation is .6 = e^(m * 2.1) which becomes .6 = e^(-243250297 * 2.1) which becomes .6 = .6.

this confirms the solution is good.

the solution is that m = -.243250297.

here's a reference on log functions that might be helpful.

https://www.purplemath.com/modules/solvelog.htm

here's another reference that might be helpful.

http://www.wtamu.edu/academic/anns/mps/math/mathlab/col_algebra/index.htm

check out tutorials 42 through 48.

any questions, write to dtheophilis@gmail.com